Problem 8-29 Completing a Master Budget [LO8-2, LO8-4, LO8-7, LO8-8, LO8-9, LO8-10]

The following data relate to the operations of Shilow Company, a wholesale distributor of consumer goods:

Current assets as of March 31:
Cash $

7,400

Accounts receivable $

19,600

Inventory $

39,000

Building and equipment, net $

126,000

Accounts payable $

23,175

Common stock $

150,000

Retained earnings $

18,825

The gross margin is 25% of sales.

Actual and budgeted sales data:

March (actual) $ 49,000
April $ 65,000
May $ 70,000
June $ 95,000
July $ 46,000

Sales are 60% for cash and 40% on credit. Credit sales are collected in the month following sale. The accounts receivable at March 31 are a result of March credit sales.

Each month’s ending inventory should equal 80% of the following month’s budgeted cost of goods sold.

One-half of a month’s inventory purchases is paid for in the month of purchase; the other half is paid for in the following month. The accounts payable at March 31 are the result of March purchases of inventory.

Monthly expenses are as follows: commissions, 12% of sales; rent, $2,200 per month; other expenses (excluding depreciation), 6% of sales. Assume that these expenses are paid monthly. Depreciation is $945 per month (includes depreciation on new assets).

Equipment costing $1,400 will be purchased for cash in April.

Management would like to maintain a minimum cash balance of at least $4,000 at the end of each month. The company has an agreement with a local bank that allows the company to borrow in increments of $1,000 at the beginning of each month, up to a total loan balance of $20,000. The interest rate on these loans is 1% per month and for simplicity we will assume that interest is not compounded. The company would, as far as it is able, repay the loan plus accumulated interest at the end of the quarter.

Required:

Using the preceding data:

1. Complete the following schedule (Expected Cash Collection):

2. Complete the following (Merchandise Purchase Budget and Schedule of Expected cash disbursments - Merchandise Purchase Budget) :

3. Complete the following cash budget:

4. Prepare an absorption costing income statement for the quarter ended June 30.
Answer #1
1) Shilow company
Schedule of Expected cash collections
April May June Quarter
Cash sales 39000 42000 57000 138000
credit sales 19,600 26000 28000 73,600
total collections 58600 68000 85000 211600
Accounts receivable = 95000*40%= 38000
2) Merchandise purchase budget
April May June Quarter
Budgeted cost of goods sold 48750 52500 71250 172500 34500
Add Desired ending inventory 42000 57000 27,600 27,600
total needs 90750 109500 98850 200100
less beginning inventory 39,000 42,000 57,000 39,000
Required purchases 51,750 67,500 41,850 161,100
cost of goods sold = 75% of sales
ending inventory = 80% of following months budgeted cost of goods sold
3) Schedule of Cash disbursements-Merchandise purhcase
April May June Quarter
March purchases 23,175 23,175
April purchases 25875 25,875 51750
May purchases 33750 33,750 67500
June purchases 20925 20925
total disbursements 49,050 59625 54675 163,350
Accounts payable june 30 = 20,925
4) Cash budget
April May June Quarter
Beginning cash balance 7,400 4,650 4,225 7,400
Add Cash collectiosn 58600 68000 85000 211600
total cas h available 66,000 72,650 89,225 219,000
less cash disbursements
for inventory 49,050 59625 54675 163,350
for expenses 13900 14800 19300 48000
for equipment 1,400 0 0 1,400
total cash disbursements 64,350 74425 73975 212,750
Excess(Deficiency)of cash 1,650 -1,775 15,250 6,250
Financing:
Borrowings 3,000 6,000 0 9,000
Repayments 0 -9,000 -9,000
interest 0 -210 -210
total financing 3,000 6,000 -9210 -210
Ending cash balance 4,650 4,225 6,040 6,040
interest = 3000*1%*3= 90
6000*1%*2= 120
210
5) income statement
Sales 230000
cost of goods sold
Beginning inventor 39,000
Add purchases 161,100
goods available for sale 200,100
ending inventory 27,600 172,500
Gross margin 57,500
Selling and administrative expense
commissions 27600
rent 6600
Depreciation (945*3) 2835
other expenses 13800 50835
net operating 6,665
interest expense 210
net income 6,455
Balance sheet
Assets
current assets
Cash 6,040
Accounts receivable 38,000
inventory 27,600
total current assets 71,640
Building And equipment ,net (126000+1400-2835) 124565
total Assets 196,205
liabilities And stockholder 's Equity
Accounts payable 20,925
total current assets 20,925
Stockholder's Equity
Capital stock 150,000
Retained earnings(18,825+6455) 25280 175,280
total liabilites & stockholders Equity 196,205